OpenStudy (anonymous):

Write the equation of the line which passes through (3, –1) and is perpendicular to the line with equation 2y – x = 1

OpenStudy (anonymous):

y=mx+B y=1/2 * x + 1/2 standard form y-y_1=m(x-x_1)+b y-(-1)=1/2(x-3) y+1=1/2 *x - 3/2 now solve into standard form and perpendicular means inverted and negative slope